OpenStudy (anonymous):

2x^2-7x+3=0 completeing the square im using the quadratic formula but i keep getting - sqare root so do i use imaginary numbers?

OpenStudy (turingtest):

the discriminant is\[b^2-4ac=(-7)^2-4(2)(3)=25>0 \]so you have made a mistake somewhere.

OpenStudy (anonymous):

There for making it -7i+_\[-7i \pm \sqrt{73}\]/4

OpenStudy (anonymous):

All over 4?

OpenStudy (turingtest):

show your work, you have made a mistake somewhere...

OpenStudy (turingtest):

\[2x^2-7x+3=0\]\[a=2\]\[b=-7\]\[c=3\]\[x={-b \pm \sqrt{b^2-4ac} \over 2a}\]plug in the numbers and tell me what you get.

OpenStudy (anonymous):

Same answer.

OpenStudy (anonymous):

-73